OpenStudy (anonymous):

how would you get the equation of the line that passes through the point (6, -2) and has a slope of -2/3?

OpenStudy (tkhunny):

If you know the Point-Slope form, you just write it down. What are the chances you know that form?

OpenStudy (anonymous):

y-y1=m(x-x1)+b where m is your slope and b is your intercept. So if you plug in your point (6,-2) and your slope m=-2/3 then this is what you'll get: y-(-2)=-2/3(x-6) y+2=-2/3(x-6) \[y+2=-\frac{ 2 }{ 3 }x+4\] \[y=-\frac{ 2 }{ 3 }x+2\] Hope that helped!

OpenStudy (tkhunny):

?? Get that "+b" out of there and you'll be right on.

OpenStudy (anonymous):

Oh whoops, my bad! Yup, the correct equation is y-y1=m(x-x1). :)

OpenStudy (anonymous):

thank you so much!!!